OpenStudy (anonymous):

Find the slope of the line passing through (3, 5) and (-7, -6).

OpenStudy (anonymous):

This problem is very similar to the last one, we use the slope formula. m=y2-y1/x2-x1. Plugging in the given coordinates we get: m=-6-5/-7-3. After simplifying we get that m=-11/-10

OpenStudy (anonymous):

11/10

OpenStudy (anonymous):

Find the slope of the line passing through (5, 5) and (-4, 5).

OpenStudy (anonymous):

Again, use the slope equation. M=y2-y1/x2-x1. Using the coordinates that are provided we get m=5-5/-4-5, this gives us 0/9, which is 0. Thus the slope is 0. This appears to be a straight line, straight line have no slope.